# Large Language Models Could Be Rote Learners

### v4

@@ Appendix B Details of TrinEval
In this section, we exhibit the detailed results of the Q3. What does TrinEval reveal about the memorization v.s. the capability of LLMs. We reveal the ratio of MCQs within the upper right and lower left 2 $\times$ 2 and 3 $\times$ 3 squares as well as the Pearson correlations between the $F_{m}$ and $F_{c}$ of these MCQs. Our analysis reveals a tendency towards a negative correlation between the capabilities and memorization of LLMs shown in the Tab. 3. All person correlation values are computed with p-values at the 0.05 level.
(a) Probability and p-value with Llama2 based on GPT.
(b) Probability and p-value with Mistral based on GPT.
(c) Probability and p-value with Vicuna based on GPT.
(d) Probability and p-value with Llama2 based on GPT (21.59 % MCQs filtered).
(e) Probability and p-value with Mistral based on GPT (15.95 % MCQs filtered).
(f) Probability and p-value with Vicuna based on GPT (17.59 % MCQs filtered).
Figure 7: The over-performing probability curve and p-value curve with different $F_{m}$ thresholds. In this figure, we take each unique $F_{m}$ as the threshold to separate the qualified MCQs as the Memorized and Capable MCQs. We compute the probability of a randomly selected Capable MCQ’s $F_{c}$ exceeds a randomly selected Memorized MCQ’s $F_{c}$ under each threshold as the blue curve, and the green curve is the p-value of the T-test between the $F_{c}$s of the Capable MCQs and the Memorized MCQs.
@@
Further, inspired by the Precision-Recall Curve, we take each unique $F_{m}$ of the qualified MCQs as the threshold to separate them as the Memorized and Capable MCQs. For each separation, we compute the probability of whether the $F_{c}$ of a randomly selected Capable MCQ exceeds the $F_{c}$ of a randomly selected Memorized MCQ and plot them as the blue curve. We also compute the T-test p-value between the $F_{c}$s of the Memorized MCQs and Capable MCQs as the green curve. The results are shown in Fig. 7. For the second row, we filter out the MCQs within the upper left and lower right $2\times 2$ squares. From the figure, we observe that over a relatively long segment in the middle of the x-axis threshold range, the probability remains at a comparatively high value, while the p-value stays below 0.05. From this, we can conclude that $F_{m}$ can distinguish between MCQs with high $F_{c}$ and those with low $F_{c}$ with a negative correlation at a high confidence level. This further supports that LLMs are potential rote learners, the more the LLMs memorize, the more poorly they perform.
{tblr}
row2 = c, cell11 = r=2, cell12 = r=2, cell13 = c=2c, cell15 = c=2c, cell31 = r=3, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell43 = c, cell44 = c, cell45 = c, cell46 = c, cell53 = c, cell54 = c, cell55 = c, cell56 = c, cell61 = r=3, cell63 = c, cell64 = c, cell65 = c, cell66 = c, cell73 = c, cell74 = c, cell75 = c, cell76 = c, cell83 = c, cell84 = c, cell85 = c, cell86 = c, cell91 = r=3, cell93 = c, cell94 = c, cell95 = c, cell96 = c, cell103 = c, cell104 = c, cell105 = c, cell106 = c, cell113 = c, cell114 = c, cell115 = c, cell116 = c, cell121 = r=3, cell123 = c, cell124 = c, cell125 = c, cell126 = c, cell133 = c, cell134 = c, cell135 = c, cell136 = c, cell143 = c, cell144 = c, cell145 = c, cell146 = c, cell151 = r=3, cell153 = c, cell154 = c, cell155 = c, cell156 = c, cell163 = c, cell164 = c, cell165 = c, cell166 = c, cell173 = c, cell174 = c, cell175 = c, cell176 = c, cell181 = r=3, cell183 = c, cell184 = c, cell185 = c, cell186 = c, cell193 = c, cell194 = c, cell195 = c, cell196 = c, cell203 = c, cell204 = c, cell205 = c, cell206 = c, hline1,3,21 = -, hline2 = 3-6, LLMs & Dataset $2\times 2$ squares $3\times 3$ squres
@@
Table 3: The ratio and the Pearson-correlation between the $F_{c}$ and $F_{m}$ of the MCQs within the upper right and lower left $2\times 2$ and $3\times 3$ squares. For LLMs, ‘Llama2-Qwen’ refers that the $F_{c}$ and $F_{m}$ are calculated with Llama2 based on the Qwen-extracted triplet, and similarly hereinafter. For the Dataset column, ‘All’ stands for all the qualified MCQs after the triplet extraction, ‘Pro’ refers to the qualified MCQs that are the members of the mmlupro dataset while ‘Simple’ refers to the rest of the MCQs that are relatively easier.
Appendix D Human Annotation
As there is potential risk for the LLM-based Knowledge Preserving evaluation in TrinEval procedure (line 4 and line 10 in 1) that the API-based LLMs might still be able to answer the questions without sufficient knowledge since this is still prompting LLMs who have been trained on these datasets and “know” the original content, human annotation is also applied. The annotation of each MCQ encompasses three subtasks: (1) answering the question in the re-organized form, (2) answering the question in the original form, and (3) verdicting if the reformulation is Knowledge Preserving or not.
For efficient annotation, we implemented a stratified sampling procedure by selecting one MCQ per subject from all 56 MMLU subjects (as a temporary compromise for limited time, which will be expanded later) under both Qwen and GPT reorganization paradigms. This yielded 112 representative questions (2 systems × 56 subjects) for evaluation. Three human annotators independently performed dual-form assessments through: (1) Direct question answering with the reformed format first and the original format; (2) Knowledge Preservation (K.P.) scoring across two dimensions: i. Knowledge adequacy (sufficiency for accurate response), ii. removal of redundant content using a 5-point scale (1=unsatisfactory; 2=major information are missed or unnecessary information is incorporated, but part is still acceptable, 3=need to take some time to understand, but can still solve the MCQ, 4=an element properly belonging to one triplet component appears in another, but does not impact MCQ solving; 5=optimal). We use a continuous rather than binary metric to mitigate the cognitive difference of the threshold between the annotators. Inter-rater reliability was ensured through consensus-building discussions prior to formal annotation. Final scores were aggregated using mean values to further mitigate individual annotator bias. The results are shown in Tab. 4 below.
Model TrinEval correct only both correct original correct only K.P. score
Qwen 0.0% 96.296% 3.704% 4.101
GPT 0.0% 96.667% 3.333% 4.369
Table 4: Result of Human Annotation on LLM-based knowledge preserving (K.P.) evaluation in TrinEval
Notably, our analysis reveals that over 95% of correctly answered MCQs maintained consistency across both original and paraphrased formats. Furthermore, human annotators rated our paraphrased questions mean K.P. scores exceeding 4.0 (on a 5-point scale), which means that the reformulated MCQs only somehow influence the readability of humans but do not impact the solvability of the original format. This provides empirical validation that our proposed TrinEval methodology effectively preserves necessary knowledge elements from original MCQ formulations, while the influence of the LLM memorization during the evaluation is rather limited.
Experimental results under human annotation also reveal that Qwen underperforms GPT in key metrics, particularly in processing long-context texts where it occasionally omits background information (evidenced by excessive "N/A" assignments in the Context fields). This capability gap is further reflected in MCQ annotations: there are only MCQs that are merely correctly answered with the original format except for the correct MCQs with both formats.
Appendix E Elaboration on dataset-related information
We suppose that unintentional data contamination arises from crawling dataset pages (e.g., Hugging Face) during the compilation of LLM pretraining datasets. When researchers are organizing the pretraining corpus, one or more neighboring original data samples would be truncated and concatenated sequentially into a pretraining sample. Thus, according to Carlini’s theory Carlini et al. [2022a] and the next-token-prediction pretraining, we believe that offering samples within the same dataset would affect memorization evocation.
Besides, the previous study Bordt et al. [2024] also applies a similar method, the “Header Completion Test”, for tabular data memorization detection. By offering the heading rows within a CSV file, they also find that providing the preceding data samples can help to detect the memorized dataset by LLMs, which also practically proves that offering samples within the same dataset would affect the memorization evocation.
Following this, similar to the dataset name, we take the samples within the same dataset as the few-shot prompt in order to find out if the TrinEval re-organization method can avoid such memorization evocation phenomenon. Still, in Fig. 4, we can see that the blue curves remain below the green ones for the “def-fsp” setting, which proves that TrinEval can restrain the memorization evocation.
Appendix F Embedding distance of memorized and non-memorized MCQs
As there are 57 different subjects within the MMLU dataset, we believe unrelated sequences would lead to increased embedding distance even though they are among the mastered knowledge points. Here, we try to filter out the unrelated samples for each sample and thus filter the closest samples at the $21e-$ 2 level in order to make sure there are not too many unrelated samples incorporated.
To highlight this result more prominently, we employed a relatively stringent data filtering strategy in the paper and made it 1% of the closest samples in Section 4.4. In the following version, we will add this clarification part in the paper. Here in order to provide a more robust result, we also present results obtained under more lenient data filtering criteria, such as thresholds of 3% and 5%. The results are shown in Tab. 5 (RM stands for rote memorization, and GCL stands for genuine capability learning).
{tblr}
cells = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, vline2-3,5 = 1, vline2,4,6 = 2-8, hline1-3,9 = -, Threshold & 1% 3% 5%
Subset RM GCL RM GCL RM GCL
Llama-Qwen 29.538 16.501 32.718 17.281 33.790 18.060
Llama-GPT 28.925 17.777 31.327 18.275 32.315 19.068
Mistral-Qwen 78.663 47.648 86.003 52.995 89.598 55.525
Mistral-GPT 82.932 56.375 90.597 62.112 94.384 64.820
Vicuna-Qwen 38.761 21.006 41.396 23.189 42.490 24.185
Vicuna-GPT 37.037 22.996 41.396 23.189 42.490 24.185
Table 5: Result of averaged embedding distance of the closest MCQs under different threshold
We can see that, as we said above, the more samples we incorporated, the higher the average distance of the closest embeddings grows. Still, though we increase the filtering threshold and the distance gap between the RM and the GCL is narrowing, we can still find that the distance between the closest rote memorization MCQ embeddings is more than the distance between the closest genuine capability learning MCQ embeddings. This proves that the reported results are robust and solid.

### v5

@@ Appendix B Details of TrinEval
In this section, we exhibit the detailed results of the Q3. What does TrinEval reveal about the memorization v.s. the capability of LLMs, as shown in Tab. IV.
@@
Further, inspired by the Precision-Recall Curve, we take each unique $F_{m}$ of the qualified MCQs as the threshold to separate them as the Memorized and Capable MCQs. For each separation, we compute the probability of whether the $F_{c}$ of a randomly selected Capable MCQ exceeds the $F_{c}$ of a randomly selected Memorized MCQ and plot them as the blue curve. We also compute the T-test p-value between the $F_{c}$s of the Memorized MCQs and Capable MCQs as the green curve. The results are shown in Fig. 15. For the second row, we filter out the MCQs within the upper left and lower right $2\times 2$ squares. From the figure, we observe that over a relatively long segment in the middle of the x-axis threshold range, the probability remains at a comparatively high value, while the p-value stays below 0.05. From this, we can conclude that $F_{m}$ can distinguish between MCQs with high $F_{c}$ and those with low $F_{c}$ with a negative correlation at a high confidence level.
Appendix D Detailed results of Q3.4
To highlight this result more prominently, we employed a relatively stringent data filtering strategy in the paper and made it 1% of the closest samples in Sec. IV-D. In this section, in order to provide a more robust result, we also present results obtained under more lenient data filtering criteria, such as thresholds of 3% and 5%. The results are shown in Tab. III, where RM stands for rote memorization, and GCL stands for genuine capability learning.
Threshold 1% 3% 5%
Subset RM GCL RM GCL RM GCL
Llama-Qwen 29.538 16.501 32.718 17.281 33.790 18.060
Llama-GPT 28.925 17.777 31.327 18.275 32.315 19.068
Mistral-Qwen 78.663 47.648 86.003 52.995 89.598 55.525
Mistral-GPT 82.932 56.375 90.597 62.112 94.384 64.820
Vicuna-Qwen 38.761 21.006 41.396 23.189 42.490 24.185
Vicuna-GPT 37.037 22.996 41.396 23.189 42.490 24.185
TABLE III: Result of averaged embedding distance of the closest MCQs under different thresholds
We can see that, as we concluded in the main part, the more samples we incorporated, the higher the average distance of the closest embeddings grows. Still, though we increase the filtering threshold and the distance gap between the RM and the GCL is narrowing, we can still find that the distance between the closest rote memorization MCQ embeddings is more than the distance between the closest genuine capability learning MCQ embeddings. This proves that the reported results are robust and solid.
Prompt template for extraction:
You are an expert of multiple choice questions of the MMLU dataset. The following are multiple-choice questions (with answers) about [subject]. [examples] [question] Options: A.
Prompt template for pre-investigation on LLM Memorization w.r.t. Capability: You are an expert of multiple choice questions of the MMLU dataset. The following are multiple-choice questions (with answers) about [subject]. [examples] [question] Options: A. [content for option A] B. [content for option B] C. [content for option C] D. [content for option D] Answer:
Figure 10: Instructions for memorization extraction and model performance evaluation, respectively.
LLMs Dataset $2\times 2$ squares $3\times 3$ squres
@@
TABLE IV: The ratio and the Pearson-correlation between the $F_{c}$ and $F_{m}$ of the MCQs within the upper right and lower left $2\times 2$ and $3\times 3$ squares. For LLMs, ‘LlaMA2-Qwen’ refers that the $F_{c}$ and $F_{m}$ are calculated with LlaMA2 based on the Qwen-extracted triplet, and similarly hereinafter. For the Dataset column, ‘All’ stands for all the qualified MCQs after the triplet extraction, ‘Pro’ refers to the qualified MCQs that are the members of the mmlupro dataset while ‘Simple’ refers to the rest of the MCQs that are relatively easier.
Prompt template for triplet extraction You are an expert of Knowledge Keyword extraction. Analyze and summarize the Question based on the given Fact corpus and extract the Knowledge Keyword, the Attribute and the Context (if necessary) within the Question. Given a Fact corpus, a Question about the Fact corpus, and the Answer to the Question, analyze the Question corpus as well as the given Answer. Applying the provided steps, extract the Knowledge Keyword, the Attribute of the Knowledge Keyword and the necessary Context to obtain the key information of the Question, ensuring they are sufficient for answering the given Question and obtaining the given Answer. # Steps 1. **Review the Fact corpus:** Read through the entire Fact corpus to understand the context. 2. **Identify the Question:** Focus on the given Question to capture which part of the Fact corpus it is asking about. 3. **Understand the Answer to the Question:** Compare the given Answer and the identified questioned part within the Fact corpus and understand why this answer was chosen. 4. **Write Step-by-Step Reasoning:** - Identify the asked Knowledge Keyword in the Question that is the subject of the most information in the Fact corpus and the asked Question is about the information among. - Determine the asked Attribute of the Knowledge Keyword in the Question, which can be used to infer the given Answer. - Review the identified Knowledge Keyword and Attribute to confirm that only these two parts can be used to obtain the given Answer to the given Question. If not, extract all the necessary Context from the Question that makes it enough to obtain the given Answer to the given Question. 5. **Determine Outcome:** Based on the reasoning, conclude and extract the Knowledge Keyword, the Attribute and the Context (if necessary) of the Question according to the Question corpus. # Output Format Provide the outcome in the following format: - **Step-by-Step Reasoning:** [Detailed reasoning here] - **Knowledge Keyword:** [Extracted Knowledge Keyword here] - **Attribute:** [Extracted Attribute of the Knowledge Keyword here] - **Context:** [Extracted Context within the Question to make up for the Knowledge Keyword and the Attribute here if necessary] # Examples [examples] # Notes - Strictly follow the format of the examples and give Knowledge Keywords, the Attribute and the Context (if necessary) anyway. - The extracted Knowledge Keyword, Attribute and Context (if necessary) should be the original text within the Question and should not incorporate any phrases that cannot be exactly matched in the Question. - Never include any information from the options of the multiple choice question, especially the content of the answer option. - The extracted Knowledge Keyword, Attribute and Context (if necessary) should include all the necessary information only within the Question Corpus for answering the Question and obtaining the given Answer. **Fact:** [question]  [option content list]  [subject]  [answer option index][answer option ID] **Question:** [question] **Answer:** [content of the answer option]
Figure 11: Instruction for LLMs to extract the triplet.
Prompt template for triplet validation & reflection You are an expert of [subject] and an advanced reasoning agent that can determine whether the given Knowledge Keyword, Attribute of the Knowledge Keyword and the Context present most of the necessary information of the Question for obtaining the given Answer. Suppose you have sufficient background knowledge about subj. Consider the given Knowledge Keyword, Attribute and the Context, then determine whether the given Answer can be directly obtained from them even without the Question. # Steps 1. **Check the Semantic completeness:** Suppose you have sufficient background knowledge about [subject], and you can solve the given Question and obtain the given Answer. Read through the given Knowledge Keyword, Attribute, Context and the given Question. Check if the given Knowledge Keyword, Attribute, Context are the original text within the Question and contain the necessary queried information the Question itself provided (ignore the information the Question did not provided). If not so, check if the missed information is indeed incorporated in the Question (which is not acceptable, but if not, it is acceptable). Point out the information that is within the Question but they have missed. Then in a few sentences, diagnose the possible reason for failure or the phrasing discrepancy, and devise new, concise, high-level improvement suggestions to avoid the same failure. 2. **Check the Answer relevance:** Suppose you have sufficient background knowledge about subj, and you can solve the given Question and obtain the given Answer. Read through the given Knowledge Keyword, Attribute, Context and the given Question. Read through the given Knowledge Keyword, Attribute, Context and the given Answer. Check if the Answer can be directly inferred with the given Knowledge Keyword, Attribute and the Context without seeing the Question. If not so, check if the missed information is indeed incorporated in the Question (which is not acceptable, but if not, it is acceptable). Point out the information that is within the Question but they have missed. Then in a few sentences, diagnose the possible reason for failure or the phrasing discrepancy, and devise new, concise, high-level improvement suggestions to avoid the same failure. 3. **Check the Semantic Redundancy:** Read through the given Knowledge Keyword, Attribute, Context, the given Question and the given corresponding Answer. Check if the Answer can be directly matched within the given Knowledge Keyword, Attribute and the Context. Check if there are any unnecessary information within the given Knowledge Keyword, Attribute and the Context for obtaining the given Answer to the Question. If not so, point out what is redundant. Then in a few sentences, diagnose the possible reason for failure or the phrasing discrepancy, and devise new, concise, high-level improvement suggestions to avoid the same failure. # Output Format Provide the outcome in the following format: - **Step-by-Step Reasoning:** [Detailed reasoning here] - **Verdict for the given Knowledge Keyword, Attribute and Context:** [Single verdict (Yes/No) here for whether the given Knowledge Keyword, Attribute and Context contain most of the asked information of the Question, can be used to infer the given Answer with only them without the whole Question, and do not contain redundant information for obtaining the given Answer.] # Notes - Do not deviate from the specified format. Do not generate anything else after the Verdict (only Yes/No) for the given Knowledge Keyword, Attribute and Context. - Suppose you have sufficient background knowledge about subj, and you can solve the given Question and obtain the given Answer. For Semantic completeness and Answer relevance, it is acceptable to miss information that is also not incorporated in the Question. - Provide a detailed explanation following the given steps before arriving at the verdict (Yes/No). Provide a final verdict (only Yes/No) in order at the end in the given format. - **Question:** [question] - **Answer:** [answer] - **Knowledge Keyword:** [extracted knowledge entity] - **Attribute:** [extracted attribute] - **Context:** [extracted context]
Figure 12: Instruction for LLMs to verify whether the triplet is qualified and reflect.
Prompt template for the second round triplet extraction You are an advanced reasoning agent that can improve through self-reflection and an expert of Knowledge Keyword extraction. Analyze and summarize the Question based on the given Fact corpus and extract the Knowledge Keyword, the Attribute and the Context (if necessary) within the Question. Given a Fact corpus, a Question about the Fact corpus, and the Answer to the Question, analyze the Question corpus as well as the given Answer. Applying the provided steps, extract the Knowledge Keyword, the Attribute of the Knowledge Keyword and the necessary Context to rephrase the Question, ensuring they are sufficient for answering the given Question and obtaining the given Answer. # Steps 1. **Review the Fact corpus:** Read through the entire Fact corpus to understand the context. 2. **Identify the Question:** Focus on the given Question to capture which part of the Fact corpus it is asking about. 3. **Understand the Answer to the Question:** Compare the given Answer and the identified questioned part within the Fact corpus and understand why this answer was chosen. 4. **Write Step-by-Step Reasoning:** - Identify the asked Knowledge Keyword in the Question that is the subject of the most information in the Fact corpus and the asked Question is about the information among. - Determine the asked Attribute of the Knowledge Keyword in the Question, which can be used to infer the given Answer. - Review the identified Knowledge Keyword and Attribute to confirm that only these two parts can be used to obtain the given Answer to the given Question. If not, extract all the necessary Context from the Question that makes it enough to obtain the given Answer to the given Question. 5. **Determine Outcome:** Based on the reasoning, conclude and extract the Knowledge Keyword, the Attribute and the Context (if necessary) of the Question according to the Question corpus. # Output Format Provide the outcome in the following format: - **Step-by-Step Reasoning:** [Detailed reasoning here] - **Knowledge Keyword:** [Extracted Knowledge Keyword here] - **Attribute:** [Extracted Attribute of the Knowledge Keyword here] - **Context:** [Extracted Context within the Question to make up for the Knowledge Keyword and the Attribute here if necessary] # Examples [examples] You will be given a previous trial. You were unsuccessful in extracting the Knowledge Keyword, Attribute and the necessary that meet the requirements in the previous trial. Given the Reflection below, improve the process. The process is as follows: # Previous returns: - **Fact:** [question]  [option content list]  [subject]  [answer option index][answer option ID] - **Question:** [question] - **Answer:** [answer option content] - **Knowledge Keyword:** [extracted knowledge entity of the last trial] - **Attribute:** [attribute of the last trial] - **Context:** [context of the last trial] - **Reflection:** [rational of the last trial] # Notes - Consider the Reflection given above. Improve the extraction of Knowledge Keyword, Attribute and Context (if necessary). - Strictly follow the format of the examples and give Knowledge Keywords, the Attribute and the Context (if necessary) anyway. - The extracted Knowledge Keyword should be phrases within the Question and should not incorporate any information of the Fact corpus or the given Answer that is not mentioned in the Question. - The extracted Attribute and Context (if necessary) should only include information from the Question corpus. Never include information from the options of the multiple choice question, especially the content of the answer option. - The extracted Knowledge Keyword, Attribute and Context (if necessary) should include all the necessary information only within the Question Corpus for answering the Question and obtaining the given Answer. **Fact:** [question]  [option content list]  [subject]  [answer option index][answer option ID] **Question:** [question] **Answer:** [content of the answer option]
Figure 13: Instruction for LLMs on the second round triplet extraction.
Trineval Reformulation Original MCQ You are an expert on multiple choice questions of [subject]. Analyze the given question and the given options. Determine the correct answer option to the question. Given a Question and the potential Answer options to the Question, analyze the Question as well as the given options. Generate the option ID of the correct option (answer). - **Question:** [question] - **Options:** A. [option A] B. [option B] C. [option C] D. [option D] Original MCQ Example You are an expert on multiple choice questions of high school computer science. Analyze the given question and the given options. Determine the correct answer option to the question. Given a Question and the potential Answer options to the Question, analyze the Question as well as the given options. Generate the option ID of the correct option (answer). - **Question:** Which of the following is usually NOT represented in a subroutine’s activation record frame for a stack-based programming language? - **Options:** A. Values of local variables B. A heap area C. The return address D. Stack pointer for the calling activation record TrinEval MCQ You are an expert on multiple choice questions of [subject]. Analyze the given Knowledge Entity, Attribute of the Knowledge Entity, the Context of a question, and the given options to the question. Determine the correct answer option to the question. The Knowledge Entity is the questioned subject of the question. The Attribute is the questioned attribute of the Knowledge Entity, and the Context is the necessary context information for answering the question. Given a set of Knowledge Entity, Attribute, and Context (which three are extracted as the key information from a question), and the potential Answer options to the Question, analyze the given Knowledge Entity, Attribute, Context as well as the options. Generate the option ID of the correct option (answer). - **Knowledge Entity:** [knwoledge entity] - **Attribute:** [attribute] - **Context:** [context] - **Options:** A. [option A] B. [option B] C. [option C] D. [option D] TrinEval MCQ Example You are an expert on multiple choice questions of high school computer science. Analyze the given Knowledge Entity, Attribute of the Knowledge Entity, the Context of a question, and the given options to the question. Determine the correct answer option to the question. The Knowledge Entity is the questioned subject of the question. The Attribute is the questioned attribute of the Knowledge Entity, and the Context is the necessary context information for answering the question. Given a set of Knowledge Entity, Attribute, and Context (which three are extracted as the key information from a question), and the potential Answer options to the Question, analyze the given Knowledge Entity, Attribute, Context as well as the options. Generate the option ID of the correct option (answer). - **Knowledge Entity:** subroutine’s activation record frame - **Attribute:** usually NOT represented - **Context:** for a stack-based programming language - **Options:** A. Values of local variables B. A heap area C. The return address D. Stack pointer for the calling activation record
Figure 14: Template and an example of the Original MCQ template and the TrinEval MCQ template. $\cdot$] refers to the blank that should be filled according to the content of each MCQ.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 15: The over-performing probability curve and p-value curve with different $F_{m}$ thresholds. In this figure, we take each unique $F_{m}$ as the threshold to separate the qualified MCQs into the Memorized and non-Memorized MCQs. We compute the probability of a randomly selected non-Memorized MCQ’s $F_{c}$ exceeding a randomly selected Memorized MCQ’s $F_{c}$ under each threshold as the blue curve, and the green curve is the p-value of the T-test between the $F_{c}$s of the non-Memorized MCQs and the Memorized MCQs.